# Deep Spectral Prior

### v1

@@ Acknowledgment
YC is funded by an AstraZeneca studentship and a Google studentship. CBS acknowledges support from the Philip Leverhulme Prize, the Royal Society Wolfson Fellowship, the EPSRC advanced career fellowship EP/V029428/1, EPSRC grants EP/S026045/1 and EP/T003553/1, EP/N014588/1, EP/T017961/1, the Wellcome Innovator Awards 215733/Z/19/Z and 221633/Z/20/Z, the European Union Horizon 2020 research and innovation programme under the Marie Skodowska-Curie grant agreement No. 777826 NoMADS, the Cantab Capital Institute for the Mathematics of Information and the Alan Turing Institute. AIAR gratefully acknowledges the support from Yau Mathematical Sciences Center, Tsinghua University.
Supplementary Material For:
@@
Appendix A Detailed Proofs of Methodological Results
This supplementary section presents the complete mathematical proofs for the results stated in the main paper. These results form the theoretical foundation of the proposed methodology, and the detailed derivations provided here aim to enhance clarity and transparency. By including these extended proofs, we offer additional insight into the logical structure and assumptions underlying our approach, which may be of interest to readers seeking a deeper understanding of the theoretical aspects of our work.
A.1 Spectral Regularisation and Implicit Bias of Deep Spectral Prior.
This section contains the detailed proofs corresponding to the results presented in Subsection 2.2 of the main paper.
Theorem 2.1 (Spectral Regularisation in DSP).  Let $f_{\theta}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$ be a convolutional neural network with parameters $\theta$, and let $z\in\mathbb{R}^{d}$ be a fixed random input. Let $y\in\mathbb{R}^{m}$ be a corrupted observation of an unknown clean image $x\in\mathbb{R}^{n}$, with a known corruption operator $\mathcal{A}:\mathbb{R}^{n}\to\mathbb{R}^{m}$. Let $\mathcal{F}:\mathbb{R}^{m}\rightarrow\mathbb{C}^{m}$ denote the discrete Fourier transform. Define the frequency loss: $\mathcal{L}_{\mathrm{DSP}}(\theta):=\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\right\|_{2}^{2}.$ Assume:
  1. 1.
    The network $f_{\theta}$ is trained using a smooth optimiser (e.g., gradient descent or Adam with small learning rate);
  2. 2.
    The network exhibits spectral bias, i.e., for training time $t$,
    $\left\|\frac{\partial}{\partial t}\mathcal{F}(f_{\theta}(z))(\omega)\right\|\text{ decreases as }\|\omega\|\text{ increases};$
  3. 3.
    The corruption noise $\eta=y-\mathcal{A}x$ is high-frequency, i.e.,
    $|\mathcal{F}(\eta)(\omega)|\text{ is large for high }\|\omega\|\text{ and small for low }\|\omega\|.$
Then the optimiser converges to $\hat{\theta}=\arg\min_{\theta}\mathcal{L}_{\mathrm{DSP}}(\theta)$ such that the reconstruction $\hat{x}=f_{\hat{\theta}}(z)$ satisfies:
$|\mathcal{F}(\mathcal{A}\hat{x})(\omega)|\approx|\mathcal{F}(y)(\omega)|\text{for low }\|\omega\|,\quad\text{and}\quad|\mathcal{F}(\mathcal{A}\hat{x})(\omega)|<|\mathcal{F}(y)(\omega)|\text{ for high }\|\omega\|.$
@@ Appendix A Detailed Proofs of Methodological Results
We analyse the behaviour of the optimiser minimising the frequency loss
$\mathcal{L}_{\mathrm{DSP}}(\theta):=\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\right\|_{2}^{2}.$
The discrete Fourier transform $\mathcal{F}$ is unitary, so the loss can be interpreted as minimising the squared error between the predicted and observed frequency spectra:
$\mathcal{L}_{\mathrm{DSP}}(\theta)=\sum_{\omega}\left|\mathcal{F}(\mathcal{A}f_{\theta}(z))(\omega)-\mathcal{F}(y)(\omega)\right|^{2}.$
By assumption (2), the network exhibits spectral bias, meaning that during optimisation, low-frequency components of the output $f_{\theta}(z)$ are learned faster than high-frequency components. Consequently, the frequencies of $\mathcal{F}(\mathcal{A}f_{\theta}(z))$ at lower $\|\omega\|$ converge more quickly to match $\mathcal{F}(y)(\omega)$ compared to those at higher $\|\omega\|$.
Simultaneously, assumption (3) asserts that the noise $\eta=y-\mathcal{A}x$ is primarily composed of high-frequency components. Hence, $\mathcal{F}(y)(\omega)$ contains noise-dominated magnitudes at high $\|\omega\|$, while the low-frequency components mostly reflect the clean signal $x$.
Now, under assumption (1), since a smooth optimiser (e.g., gradient descent or Adam with a small learning rate) is used, the optimisation proceeds gradually. Because of spectral bias, the model first fits the low-frequency components of $\mathcal{F}(y)$, which are likely to be signal-dominated, before fitting the high-frequency components, which are noise-dominated.
If optimisation is stopped early (or the optimiser naturally converges slowly on high frequencies), the final network output $\hat{x}=f_{\hat{\theta}}(z)$ will match $y$ in the frequency domain primarily at low frequencies. That is,
$|\mathcal{F}(\mathcal{A}\hat{x})(\omega)|\approx|\mathcal{F}(y)(\omega)|\quad\text{for low }\|\omega\|,$
while for high frequencies, due to the slow convergence and absence of strong signal structure,
$|\mathcal{F}(\mathcal{A}\hat{x})(\omega)|<|\mathcal{F}(y)(\omega)|.$
Thus, the spectral bias acts as a regulariser, suppressing overfitting to high-frequency noise. This completes the proof. ∎
Proposition 2.1 (Gradient Dynamics in Frequency Space).  Consider the frequency-domain loss $\mathcal{L}_{\mathrm{DSP}}(\theta):=\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\right\|_{2}^{2}.$ Then, when the parameters $\theta$ are updated by a smooth optimisation algorithm with small learning rate, the iterates $\{\theta_{t}\}$ yield reconstructions $u_{t}:=\mathcal{A}f_{\theta_{t}}(z)$ such that
$\mathcal{F}(u_{t})\to\mathcal{F}(y)\quad\text{in }\mathbb{C}^{m}.$
@@
We are given the frequency-domain loss function
$\mathcal{L}_{\mathrm{DSP}}(\theta):=\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\right\|_{2}^{2},$
and we assume that the parameters $\theta$ are updated by a smooth optimisation algorithm (e.g., gradient descent or Adam) with a sufficiently small learning rate.
Under this setting, standard results from optimisation imply that the iterates $\{\theta_{t}\}$ will converge to a minimiser $\hat{\theta}$ of the loss, or approach such a minimiser asymptotically in the limit of infinite iterations. Define $u_{t}:=\mathcal{A}f_{\theta_{t}}(z)$, the reconstruction at iteration $t$. Since the loss is defined in terms of the $\ell_{2}$-distance in the frequency domain, the convergence of $\theta_{t}$ implies:
$\mathcal{F}(u_{t})\to\mathcal{F}(y)\quad\text{in }\mathbb{C}^{m}.$
Furthermore, suppose that the network $f_{\theta}$ exhibits spectral bias — that is, during training, the network tends to fit low-frequency components before high-frequency ones. This behaviour has been empirically observed in various neural architectures, particularly convolutional networks.
As a consequence, the low-frequency components of the signal $\mathcal{F}(u_{t})$ will converge more rapidly to those of $\mathcal{F}(y)$, while the high-frequency components will evolve more slowly. More precisely, if $\omega\in\mathbb{R}^{m}$ denotes a frequency, then:
$\left|\mathcal{F}(u_{t})(\omega)-\mathcal{F}(y)(\omega)\right|\text{ decays faster for small }\|\omega\|\text{ than for large }\|\omega\|.$
This completes the proof. ∎
A.2 Optimisation Properties of Deep Spectral Prior
The following proofs support the theoretical results presented in Subsection 2.3 of the main text.
Theorem 2.2 (Smoothness and Local Stability of our Frequency Loss).  Let $f_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{n}$ be a neural network that is continuously differentiable with respect to parameters $\theta\in\mathbb{R}^{p}$. Let $\mathcal{A}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be a fixed linear corruption operator, and let $\mathcal{F}:\mathbb{R}^{m}\to\mathbb{C}^{m}$ denote the discrete Fourier transform (DFT). For a fixed input $z\in\mathbb{R}^{d}$ and observed corrupted image $y\in\mathbb{R}^{m}$, define the DSP frequency loss as: $\mathcal{L}_{\mathrm{DSP}}(\theta):=\left\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\right\|_{2}^{2}.$ It follows that:
  1. 1.
    $\mathcal{L}_{\mathrm{freq}}$ is continuously differentiable with respect to $\theta$;
  2. 2.
    $\mathcal{L}_{\mathrm{freq}}$ is locally Lipschitz in $\theta$;
  3. 3.
    Any local minimum $\theta^{*}$ satisfies the first-order optimality condition $\nabla_{\theta}\mathcal{L}_{\mathrm{freq}}(\theta^{*})=0$;
  4. 4.
    In a neighborhood of any stationary point $\theta^{*}$, gradient-based optimisation (e.g., gradient descent or Adam with sufficiently small step size) yields descent in the frequency domain,
    $\mathcal{L}_{\mathrm{freq}}(\theta_{t+1})\leq\mathcal{L}_{\mathrm{freq}}(\theta_{t}),$
    for all $t$ sufficiently large.

### v2

@@ Acknowledgment
YC is funded by an AstraZeneca studentship and a Google studentship. XZ thanks Qiuzhen College, Tsinghua University, for supporting this research. CBS acknowledges support from the Philip Leverhulme Prize, the Royal Society Wolfson Fellowship, the EPSRC advanced career fellowship EP/V029428/1, EPSRC grants EP/S026045/1 and EP/T003553/1, EP/N014588/1, EP/T017961/1, the Wellcome Innovator Awards 215733/Z/19/Z and 221633/Z/20/Z, the European Union Horizon 2020 research and innovation programme under the Marie Skodowska-Curie grant agreement No. 777826 NoMADS, the Cantab Capital Institute for the Mathematics of Information and the Alan Turing Institute. AIAR gratefully acknowledges the support from Yau Mathematical Sciences Center, Tsinghua University. This work is also supported by the Tsinghua University Dushi Program.
@@
Appendix A Supplementary Theoretical Proofs
This supplementary section presents the complete mathematical proofs for the results stated in the main paper. These results form the theoretical foundation of the proposed methodology, and the detailed derivations provided here aim to enhance clarity. By including these extended proofs, we offer additional insight into the logical structure and assumptions underlying our approach, which may be of interest to readers seeking a deeper understanding of the theoretical aspects of our work.
Corollary 2.1 (Equivalence of Deep Spectral Prior Loss and Pixel-Space Error).  Let $x,\hat{x}\in\mathbb{R}^{m}$, and let $\mathcal{F}:\mathbb{R}^{m}\to\mathbb{C}^{m}$ denote the unitary discrete Fourier transform. Then
$\|x-\hat{x}\|_{2}^{2}=\|\mathcal{F}(x)-\mathcal{F}(\hat{x})\|_{2}^{2}.$
In particular, for any neural network output $f_{\theta}(z)$ and observation $y$, $\mathcal{L}_{\mathrm{DSP}}(\theta)=\frac{1}{2}\big\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\big\|_{2}^{2}$ is exactly equal to the pixel-domain reconstruction error $\frac{1}{2}\|\mathcal{A}f_{\theta}(z)-y\|_{2}^{2}$. Thus, DSP does not alter the objective value but reformulates the optimisation in the spectral domain.
@@ Appendix A Supplementary Theoretical Proofs
Let $x,\hat{x}\in\mathbb{R}^{H\times W}$ be images of spatial size $H\times W$. We use the unitary 2D discrete Fourier transform $\mathcal{F}:\mathbb{R}^{H\times W}\to\mathbb{C}^{H\times W}$. This result follows directly from Parseval’s theorem as follows:
$\|\mathcal{F}(x)-\mathcal{F}(\hat{x})\|_{2}^{2}$ $=\sum_{p=1}^{H}\sum_{q=1}^{W}\left|\mathcal{F}(x)_{p,q}-\mathcal{F}(\hat{x})_{p,q}\right|^{2}.$ (11)
Using the unitary 2D DFT
$\mathcal{F}(x)_{p,q}=\frac{1}{\sqrt{HW}}\sum_{m=1}^{H}\sum_{n=1}^{W}x_{m,n}\,e^{-2\pi i\left(\frac{pm}{H}+\frac{qn}{W}\right)},$
(and analogously for $\hat{x}$), the difference becomes
$\|\mathcal{F}(x)-\mathcal{F}(\hat{x})\|_{2}^{2}$
$=\frac{1}{HW}\sum_{p,q}\left|\sum_{m,n}(x_{m,n}-\hat{x}_{m,n})\,e^{-2\pi i\left(\frac{pm}{H}+\frac{qn}{W}\right)}\right|^{2}$
$=\frac{1}{HW}\sum_{p,q}\;\sum_{m,n}\sum_{k,\ell}(x_{m,n}-\hat{x}_{m,n})(x_{k,\ell}-\hat{x}_{k,\ell})^{*}$
$\qquad\qquad\times e^{-2\pi i\left(\frac{pm}{H}+\frac{qn}{W}\right)}e^{2\pi i\left(\frac{pk}{H}+\frac{q\ell}{W}\right)}.$ (12)
Using orthogonality identities
$\frac{1}{H}\sum_{p=1}^{H}e^{-2\pi i\,p(m-k)/H}=\delta_{m,k},$ (13)
$\frac{1}{W}\sum_{q=1}^{W}e^{-2\pi i\,q(n-\ell)/W}=\delta_{n,\ell},$ (14)
all cross-terms collapse, and we obtain
$\|\mathcal{F}(x)-\mathcal{F}(\hat{x})\|_{2}^{2}$ $=\sum_{m,n}|x_{m,n}-\hat{x}_{m,n}|^{2}$ (15)
$=\|x-\hat{x}\|_{2}^{2}.$ (16)
Thus the pixel-space and Fourier-domain errors are identical. ∎
Theorem 2.1 (Complex Descent Law: Distinct Optimisation Paths of DIP and DSP).  Consider the frequency-domain loss
$\mathcal{L}_{\mathrm{DSP}}(\theta)$ $=\frac{1}{2}\big\|\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\big\|_{2}^{2}$ (17)
$=\frac{1}{2}\big(\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\big)^{\dagger}\big(\mathcal{F}(\mathcal{A}f_{\theta}(z))-\mathcal{F}(y)\big).$
where $\mathcal{F}$ denotes the discrete Fourier transform and the Hermitian transpose. The optimisation dynamics under gradient descent differ fundamentally between real-valued DIP and complex-valued DSP:
  • DIP Optimisation: for $\theta\in\mathbb{R}^{p}$,
    $\frac{d\theta}{dt}=-\alpha\nabla_{\theta}\mathcal{L}_{\mathrm{DIP}}(\theta)$ (18)
    $=-\alpha\,\nabla_{\theta}\big(\mathcal{A}f_{\theta}(z)-y\big)^{T}\!\cdot\!\big(\mathcal{A}f_{\theta}(z)-y\big),$
    $\text{where}\,\,\mathcal{L}_{\mathrm{DIP}}(\theta)=\frac{1}{2}\big\|\mathcal{A}f_{\theta}(z)-y\big\|_{2}^{2}.$
  • DSP Optimisation: for $\tilde{\theta}=a+ib\in\mathbb{C}^{p}$,
[…]
    $=-\alpha\,\nabla_{\tilde{\theta}^{*}}\Big[\big(\mathcal{F}(\mathcal{A}f_{\tilde{\theta}}(z))-\mathcal{F}(y)\big)^{\dagger}\!\cdot\!\big(\mathcal{F}(\mathcal{A}f_{\tilde{\theta}}(z))-\mathcal{F}(y)\big)\Big].$
    where $\nabla_{\tilde{\theta}^{*}}$ denotes the Wirtinger gradient with respect to the complex conjugate variable.
@@
For gradient descent method with $\theta\in\mathbb{R}^{p}$, the parameter $\theta$ updates as
$\theta^{(t+1)}=\theta^{(t)}-\alpha\frac{\partial\mathcal{L}}{\partial\theta}(\theta^{(t)}).$ (20)
By extending $t$ as a continuous variable, the finite differentiation $\frac{\theta^{(t+1)}-\theta^{(t)}}{(t+1)-t}\simeq\frac{d\theta^{(t)}}{dt}+O(\alpha^{2})$. And as $\alpha\rightarrow 0$, the $O(\alpha^{2})$ could be ignored, and the optimisation path is approximated as
$\frac{d\theta}{dt}=-\alpha\frac{\partial\mathcal{L}}{\partial\theta}(\theta^{(t)}).$ (21)
To strictly prove this ODE approximation hold, we assume
  • local $M$-Lipshitz condition of $\mathcal{L}$:
    $\|\mathcal{L}(u)-\mathcal{L}(v)\|\leq M\|u-v\|.$
  • local $K$-Lipshitz condition of $\partial_{\theta}\mathcal{L}$ :
    $\|\frac{\partial\mathcal{L}}{\partial\theta}(u)-\frac{\partial\mathcal{L}}{\partial\theta}(v)\|\leq K\|u-v\|.$
Under these assumptions, the ODE’s one–step update satisfies
$\hat{\theta}^{(t+1)}-\theta^{(t)}=-\int_{0}^{1}\alpha\partial_{\theta}\mathcal{L}(\theta^{(u+t)})du$
$=-\alpha\partial_{\theta}\mathcal{L}(\theta^{(t)})-\int_{0}^{1}\alpha[\partial_{\theta}\mathcal{L}(\theta^{(u+t)})-\partial_{\theta}\mathcal{L}(\theta^{(t)})]du$
$=\frac{d\theta^{(t)}}{dt}-\int_{0}^{1}\alpha[\partial_{\theta}\mathcal{L}(\theta^{(u+t)})-\partial_{\theta}\mathcal{L}(\theta^{(t)})]du$ (22)
And the discrepancy between finite differentiation and derivative is
$\|\theta^{(t+1)}-\theta^{(t)}-\frac{d\theta^{(t)}}{dt}\|=\|\hat{\theta}^{(t+1)}-\theta^{(t)}+\alpha\partial_{\theta}\mathcal{L}(\theta^{(t)})\|$
$\leq\alpha\int_{0}^{1}\|\partial_{\theta}\mathcal{L}(\theta^{(u+t)})-\partial_{\theta}\mathcal{L}(\theta^{(t)})\|du$
$\leq\alpha K\int_{0}^{1}\|\theta^{(u+t)}-\theta^{(t)}\|du$ (23)
$\leq\alpha^{2}K\int_{0}^{1}\int_{0}^{u}\|\partial_{\theta}\mathcal{L}(\theta^{(s+t)})\|dsdu\leq\frac{1}{2}\alpha^{2}KM\simeq O(\alpha^{2})$
Thus, the ODE approximation (21) holds. Substituting the DIP loss into this expression yields to (18).
Now for our Deep Spectral Prior (DSP), the discrete gradient descent updates for DSP are
[…]
$\tilde{\theta}^{(t+1)}-\tilde{\theta}^{(t)}=(a^{(t+1)}-a^{(t)})+i(b^{(t+1)}-b^{(t)})$
[…]
This shows that one needs to do gradient descent for both real part $a$ and imaginary part $b$ simultaneously in one step.
One can still express the procedure as ODEs, but here we should assume the $M$-Lipshitz condition for $\mathcal{L}_{\mathrm{DSP}}$ and $K$-Lipshitz condition for both $\nabla_{a}\mathcal{L}_{\mathrm{DSP}}$ and $\nabla_{b}\mathcal{L}_{\mathrm{DSP}}$ (i.e. $\nabla_{\tilde{\theta}}\mathcal{L}_{\mathrm{DSP}}$ and $\nabla_{\tilde{\theta}^{*}}\mathcal{L}_{\mathrm{DSP}}$ ) instead. Then
$\hat{a}^{(t+1)}-a^{(t)}=-\int_{0}^{1}\alpha\partial_{a}\mathcal{L}(\tilde{\theta}^{(u+t)})du$
$=-\alpha\partial_{a}\mathcal{L}(\tilde{\theta}^{(t)})-\int_{0}^{1}\alpha[\partial_{a}\mathcal{L}(\tilde{\theta}^{(u+t)})-\partial_{a}\mathcal{L}(\tilde{\theta}^{(t)})]du$
$=\frac{da^{(t)}}{dt}-\int_{0}^{1}\alpha[\partial_{\theta}\mathcal{L}(\tilde{\theta}^{(u+t)})-\partial_{\theta}\mathcal{L}(\tilde{\theta}^{(t)})]du$ (25)
$\|\hat{a}^{(t+1)}-a^{(t)}-\frac{da}{dt}\|=\|\hat{a}^{(t+1)}-a^{(t)}+\alpha\partial_{a}\mathcal{L}(\tilde{\theta}^{(t)})\|$
$\leq\alpha\int_{0}^{1}\|\partial_{a}\mathcal{L}(\tilde{\theta}^{(u+t)})-\partial_{a}\mathcal{L}(\tilde{\theta}^{(t)})\|du$
$\leq\alpha K\int_{0}^{1}\|\tilde{\theta}^{(u+t)}-\tilde{\theta}^{(t)}\|du$
$\leq\alpha^{2}K\int_{0}^{1}\int_{0}^{u}\|(\partial_{a}+i\partial_{b})\mathcal{L}(\theta^{(s+t)})\|dsdu$
$\leq\frac{\sqrt{2}}{2}\alpha^{2}KM\simeq O(\alpha^{2})$ (26)
The $b$ part is similar. That’s why we denote GD updates of both $a$ and $b$ as ODE.
[…]
So add the real part updates together with $i$ times that of imaginary part, we obtain
$)\displaystyle\frac{d\tilde{\theta}}{dt}=\frac{da}{dt}+i\frac{db}{dt}=-\alpha\,(\nabla_{a}\mathcal{L}_{\mathrm{DSP}}(\tilde{\theta})+i\nabla_{b}\mathcal{L}_{\mathrm{DSP}}(\tilde{\theta}$))
[…]
$=-\alpha\,\nabla_{\tilde{\theta}^{*}}\Big[\big(\mathcal{F}(\mathcal{A}f_{\tilde{\theta}}(z))-\mathcal{F}(y)\big)^{\dagger}\!\cdot\!\big(\mathcal{F}(\mathcal{A}f_{\tilde{\theta}}(z))-\mathcal{F}(y)\big)\Big].$
Theorem 2.2 (Robust Stopping Property of DSP).  For a DSP optimisation trajectory, the NTK approximation takes the form
$\mathcal{F}(\mathcal{A}f_{\tilde{\theta}}(z))=\mathcal{F}(\mathcal{A}f_{\tilde{\theta}_{0}}(z))+\tilde{\Phi}(z)\tilde{u}$, (28)
[…]
where $\alpha$ is the gradient descent step size. The evolution of the discrepancy along gradient descent follows a linear ordinary differential equation (ODE) governed by the Hermitian operator $\tilde{\Phi}\tilde{\Phi}^{\dagger}$:
[…]
The ODE admits the closed-form solution
$\mathcal{F}(\mathcal{A}f_{\tilde{\theta}^{(t)}}(z))=\sum_{k=1}^{N}(1-e^{-\alpha\mu_{k}t})\langle\tilde{y},\tilde{\psi}_{k}\rangle\tilde{\psi}_{k},$ (31)
where $\tilde{\psi}_{k}$ and $\mu_{k}$ satisfy the eigendecomposition $\tilde{\Phi}\tilde{\Phi}^{\dagger}=\sum_{k=1}^{N}\mu_{k}\tilde{\psi}_{k}\tilde{\psi}_{k}^{\dagger}$. The mean-squared error (MSE) along the optimisation path of DSP evolves as
$MSE_{t}=\mathbb{E}\|\mathcal{F}(\mathcal{A}f_{\tilde{\theta}^{(t)}}(z))-\tilde{x}\|^{2}$ (32)
$=\sum_{k=1}^{N}\!\Big[(1-e^{-\alpha\mu_{k}t})^{2}(\mathbb{E}|\langle\tilde{\epsilon},\tilde{\psi}_{k}\rangle|^{2}+|\langle\tilde{\mathcal{A}x},\tilde{\psi}_{k}\rangle|^{2})$
$-2(1-e^{-\alpha\mu_{k}t})Re\big\{\langle\tilde{\mathcal{A}x},\tilde{\psi}_{k}\rangle^{*}\langle\tilde{x},\tilde{\psi}_{k}\rangle\big\}+|\langle\tilde{x},\tilde{\psi}_{k}\rangle|^{2}\Big]\!.$
where $\tilde{x}=\mathcal{F}(x)$, $\tilde{\mathcal{A}x}=\mathcal{F}(\mathcal{A}x)$ and noise $\tilde{\epsilon}=\mathcal{F}(\epsilon)$. The optimal stopping time for each mode $k$ is then
$t_{k}^{*}$ $=\frac{1}{\alpha\mu_{k}}\ln\!\left(\rho_{k}\right),$ (33)
$\rho_{k}$ $=\frac{\mathbb{E}|\langle\tilde{\epsilon},\tilde{\psi}_{k}\rangle|^{2}+|\langle\tilde{\mathcal{A}x},\tilde{\psi}_{k}\rangle|^{2}}{\mathbb{E}|\langle\tilde{\epsilon},\tilde{\psi}_{k}\rangle|^{2}+|\langle\tilde{\mathcal{A}x},\tilde{\psi}_{k}\rangle|^{2}-Re\{\langle\tilde{\mathcal{A}x},\tilde{\psi}_{k}\rangle^{*}\langle\tilde{x},\tilde{\psi}_{k}\rangle\}}.$
Different eigenvalues $\mu_{k}$ induce distinct convergence speeds, endowing DSP with an inherent robustness to stopping time. (See proof in the Supplementary Material)